When the following oxidation-reduction equation representing a reaction that takes
place in BASIC solution is correctly balanced using the smallest possible whole number
coefficients, the coefficient before the Al is:
Al + NO2
- = Al(OH)4
- + NH3
1
Expert's answer
2020-06-19T07:27:18-0400

Solution.

"Al^0 + NO2^- = Al(OH)4^- + NH3"

"Al^0 + 4OH^- = Al(OH)4^- + 3e"

"NO2^- + 5H2O + 6e = NH3 + 7OH^-"

"2Al^0 + NO2^- + OH^- + 5H2O = 2Al(OH)4^- + NH3"

Answer:

"2Al^0 + NO2^- + OH^- + 5H2O = 2Al(OH)4^- + NH3"